a package is drop from the drone which is raising with velocity 0.5m/sec.find the distance between package and drone after 2.5sec?
1
Expert's answer
2020-09-28T08:45:51-0400


When a package is dropped from drone its initial velocity will be

"u_p = -0.5 \;m/sec"

Taking upward vertical direction as negative.

"S_2 = u_pt + \u00bd at^2"

"S_2 = -0.5\times 2.5 + ½ \times 9.8 \times (2.5)^2"

"S_2 = -1.25 + 30.625 = 29.375\; m"

"S_1 = u_dt + \u00bdat^2"

a – acceleration of the drone (a = 0)

"S_1 = u_dt"

"S_1 = -0.5 \times 2.5 = -1.25 \;m"

Negative sign indicates upward direction of displacement.

"S = S_1 + S_2"

"S = 1.25 + 29.375 = 30.625\; m"

Answer: 30.625 m